Nitrous oxide (N2O) is used as an anesthetic (laughing gas) and in aerosol cans to produce whipped cream. It is also a potent greenhouse gas. The decomposition of N2O is first order. It decomposes slowly to N2 and O2:

How many half-lives will it take for the concentration of the N2O to reach 25.00 % of its original concentration?
Answer #1

Solution:-

N / N0 = (1/2)n

Where, N = final concentration

N0 = original concentration

n = number of half-lives.

Let, N0 = 100%

then, N = 25%

Thus,

(25 / 100) = (1/2)n

(1/4) = (1/2)n

(1/2)2 = (1/2)n

n = 2

Hence, the number of half-lives required for the concentration of the N2O to reach 25% of it's original concentration is = 2

----------------------------------------------------------------------